Hello everyone! I need help on the following problems. I'm answering 6 exercises but skipped the following questions from different exercises since I am very confused. I would be very happy to receive a response.
1. In how many ways can 5 people be seated in a car, two in the front seat and 3 in the rear seat, if only 3 of the 5 can drive and others cannot drive?
2. An engineering firm wishes to hire 3 plumbers, 2 electricians and 2 machine operators. How many different choices can be made if 6 plumbers, 5 electricians, and 4 machine operators are available?
3. A team consists of 10 men and 20 women of which half the men and the women have brown eyes. Find the probability that a person chosen at random is a man or has brown eyes.
4. Find: P(Z > 2.75)
5. The lifetime of flashlight batteries is normally distributed with a mean of 1.5 and standard deviation of 0.5 hours. What is the probability that a battery selected at random will have a lifetime between 2 to 3 hours?

Thank you so much!
Kind regards,
DLlamera

Answer by Boreal(15235)

2. 6C3*5C2*4C2 =20*10*6=1200.
3.===M======W====total
BE== 5======10=====15
not==5======10=====15
total==10====20=====30
Probability of a man is 1/3+probability has brown eyes=1/2- probability of both =1/6
2/3. What doesn't work is not a man and doesn't have brown eyes. That is 1/3, or 10/30, lower right, and what remains is 2/3.
4.P z>2.75 is small by empirical rule, between .13% and 2.5%, and it is 0.003, or 0.3%
5.Between 2 and 3 hours is between 1 and 3 sd s above the mean.
Probability =0.1573.